a) Let D be a nonempty subset of R, bounded above. Is M = sup(D) a limit point of D? Explain.

b) Let the function g : R → R be bounded. Define f(x) = 1 + 4x + x^2(g(x)). Prove that

f(0) = 1 and f'(0) = 4. (You may not assume g is differentiable.)

c) Suppose that the function f : R → R is differentiable and that there is a bounded sequence {xn} with xn doesnt= xm if n doesnt= m, such that f(xn) = 0 for every index n. Show that there is a point x0 at which f(x0) = 0 and f'(x0) = 0. (Hint: Use Bolzano-Weierstrass Theorem.)


1
Expert's answer
2020-10-14T15:15:28-0400

a)

let us an example;

D = {1,2,3,4}

D is non empty

D is bounded above

Supremum of D = 4

'4' is not a limit point of D

for "\in" = 1/2

4 "\in" (4 - 1/2, 4+ 1/2)"\notin" 4

therefore is not limit point

supremum of D is not a limit point for D

b)

f(x) = 1 +4x + x 2 (gx)

if x = 0

f(x) = 1 + 0 +0

f(x) = 1

therefore g is bounded

so g * 0 = 0

f(0) = 1

lim ( f(x) - f(0))/(x - 0)

x "\to" 0

= (1 + 4x + x2 - 1)/ x

= lim 4 + x

x"\to"0

= 4

thus lim (f(x) - f(0))/(x - 0) = 4 which is finite value

x"\to" 0

so we have f'(0) = 2

c)

Let f: IR"\to" IR be a function, then lim f(x) = "\rho" "\isin" IR if and only if for any sequence (fn) in IR with n"\to"a, f(m/ "\to" "\rho" n"\to" a

Given f : IR"\to"IR is differentiable and x(n) "\subset" IR with x(n) "\notin" x(m) if n "\notin" m such that f(xn) = 0 "\forall" n "\isin" IN and xn is bounded sequence. Hence by Bolzano-Weierstrass theorem, "\exist" a subsequence (xnR) of (xn) such that (xnR) convergent and converges to xo "\in" IR

Now f'(x0) = lim (f(x) - f(x0))/(x - x0) exist

x"\to" x0


we know xnR"\to" x0 and hence;


lim (f(xnR) - f(x0))/ (xnR - x0)

R"\to" "\empty"

lim (0 - 0)/ (xnR - x0)

R"\to" 0

= 0

Hence f'(x0) = lim (f(x) - f(x0))/ (x - x0)

x"\to"x0

= lim (f(xnR) - f(x0)) / (xnR - x0)

R"\to\empty"

= 0

Therefore f'(x0) = 0